### v2

# A Backbone for Long-Horizon Robot Task Understanding

Abstract
End-to-end robot learning, particularly for long-horizon tasks, often results in unpredictable outcomes and poor generalization. To address these challenges, we propose a novel Therblig-based Backbone Framework (TBBF) to enhance robot task understanding and transferability. This framework uses therbligs (basic action elements) as the backbone to decompose high-level robot tasks into elemental robot configurations, which are then integrated with current foundation models to improve task understanding. The approach consists of two stages: offline training and online testing. During the offline training stage, we developed the Meta-RGate SynerFusion (MGSF) network for accurate therblig segmentation across various tasks. In the online testing stage, after a one-shot demonstration of a new task is collected, our MGSF network extracts high-level knowledge, which is then encoded into the image using Action Registration (ActionREG). Additionally, Large Language Model (LLM)-Alignment Policy for Visual Correction (LAP-VC) is employed to ensure precise action execution, facilitating trajectory transfer in novel robot scenarios. Experimental results validate these methods, achieving 94.37% recall in therblig segmentation and success rates of 94.4% and 80% in real-world online robot testing for simple and complex scenarios, respectively. Supplementary material is available at: https://sites.google.com/view/therbligsbasedbackbone/home
I Introduction
Understanding robot tasks encompasses several key stages: sensing the environment, recognizing task-related objects, making decisions, and planning trajectories. Recently, data-driven methods, especially deep learning algorithms, have greatly advanced the field of robotics. While deep learning excels in object recognition and reinforcement learning aids in trajectory planning, these models often struggle to generalize beyond trained scenarios, especially in long-horizon tasks. Thus, improving generalization is crucial for adapting to diverse, dynamic, real-world situations effectively.
A major challenge is data efficiency, as current systems require large, resource-intensive datasets [1, 2, 3]. Models trained for long-horizon tasks excel in simple tasks such as pick&place but struggle with more complex tasks like liquid pouring, which involve multiple steps like grasping, pouring, and releasing. The narrow focus on simple tasks limits these models’ applicability in diverse and complex scenarios. Additionally, end-to-end systems are difficult to explain, analyze, and improve. These systems [4, 5, 6] can learn directly from sensory inputs and discover complex feature representations, eliminating the need for manual feature engineering. However, they are hard to interpret and struggle to transfer skills to new tasks.
To address these challenges, we propose a structured and modular backbone framework. In our context, a backbone refers to a structured framework that decomposes high-level tasks into fundamental units, each representing a specific action or sequence of actions, facilitating better task understanding and transferability [7]. Complex tasks involve rich contact interactions (e.g., multiple points of force application) and operations in dynamic or cluttered environments, requiring the integration of diverse skills and sensor modalities (e.g., vision, force) [8]. Long-horizon tasks require extended sequences of actions, typically involving more than 10 individual steps [9], or over a prolonged period (e.g., several minutes) to achieve a specific goal.
Figure 1: Illustration of the proposed robot task understanding system. By extracting the key backbone of complex tasks and integrating contextual information from one-shot demonstrations, this system efficiently reasons about task-related objects and actions, and generates new trajectories in novel scenarios.
Many models depend heavily on pre-training in specific scenarios [10, 11, 12, 13, 14, 15], limiting their ability to generalize to new environments. Effective scenario generalization is essential for adapting to the complexity and variability of real-world situations [15]. Furthermore, task generalization is crucial, as many systems require extensive examples to learn new tasks, which is impractical in dynamic environments with frequent task variations. Handling a diverse range of tasks is essential for practical robotics. However, existing models often excel only in narrow tasks and struggle with broader activities, reducing their real-world utility. Table I compares various robot learning systems in key capacities, including data efficiency, task complexity, explanation, diversity, generalization, success rate, pre-training scenario requirements, scenario complexity, and multi-modal fusion capability.
TABLE I: Comparison of Capacities of Different Robot Systems
Capacity / System Ours Q-attention [12] PerAct [10] Diffuser [13] MimicPlay [14] GLiDE [15] Coarse to Fine [11]
Data Efficiency Highly Efficient Moderately Moderately Efficient Inefficient Highly Efficient Highly Efficient Moderately Efficient
Task Complexity Long-Horizon Short-Horizon Short-Horizon Short-Horizon Long-horizon Short-Horizon Short-Horizon
Task Explanation High Clarity Low Clarity Medium Clarity Low Clarity Medium Clarity High Clarity Medium Clarity
Task Diversity Wide Domain Narrow Domain Narrow Domain Narrow Domain Wide Domain Narrow Domain Wide Domain
Task Generalisation One-shot Few-shot 53 shots 10k shots 20&40 shots Few-shot One-shot
Success Rate 94.4% None *65% *54.4% *55% *99.6% *70%
Pre-train Scenarios Non-essential Required Required Required Required Required Required
Scenario Complexity Complex Moderate Moderate Moderate Moderate Simple Simple
Multi-modal Fusion Action, Vision, LLM Vision Vision, LLM Action, Vision Action, Vision Action, Vision, LLM Action, Vision
*The results are from the original presented papers.
To overcome these limitations, we propose a structured and modular backbone framework. This approach enhances the adaptability and efficiency of robot learning systems, improving generalization across various tasks and scenarios while providing a clearer structure for task execution. Our framework (Fig. 1) decomposes robot tasks into two sub-levels: robot trajectory (internal information) and environmental context (external information). The robot trajectory covers the entire motion during the task, while the environmental context includes all objects involved.
By decomposing the robot trajectory, we identify elemental motions, termed ’Therbligs’ (details in Section 2), which serve as backbone elements for understanding the task trajectory and context. Decomposing the environmental context into lower-level elements reveals objects and their attributes, such as mask, position, orientation, and task relevance. As shown in Table I, our approach offers high data efficiency, achieving superior performance with smaller datasets, which is crucial for practical applications. The main contributions of this work are as follows:
  • Therblig-Based Backbone Framework (TBBF): A framework that decomposes complex tasks into fundamental units, enhancing generalization and data efficiency in long-horizon tasks.
  • Meta-RGate SynerFusion Network (MGSF): A network for accurate therblig segmentation during offline training, improving understanding of sequential actions.
  • Action Registration (ActionREG): Integrates therbligs with object configurations for precise action registration and robust task execution.
  • LLM-Alignment Policy for Visual Correction (LAP-VC): Improves tolerance to errors and reduces dependency on highly accurate demonstrations.
We conducted offline and online tests to evaluate the robustness and generalization of our framework. During offline training, the MGSF network segmented therbligs across six tasks: tool-pick&place, crossbeam-cutting, bricks-gluing, tissue-sweeping, surface-wiping, and cup-pouring. In online testing, we collected one-shot demonstrations for five new tasks: board-rolling, foamblock-flipping, plate-scrubbing, spoon-tilting, and paper-stamping.
II Related research
Various intelligent robot systems achieve high accuracy in specific tasks, such as cable routing [16], cloth manipulation [17], and fruit grasping [18]. However, they often struggle to generalize across different tasks. Recently, efforts have been made to develop systems that can handle a variety of tasks [14, 11, 13]. However, these systems typically require large datasets or simple scenarios for generalization and lack a clear backbone for task understanding. To overcome these limitations, we propose using therbligs as the backbone for better task comprehension. Therbligs, which consist of 18 basic motions, provide a systematic way to describe and analyze detailed actions in various tasks [19]. Initially developed for studying human movements, we apply these concepts to visually and conceptually break down the robot’s trajectory during learning and task execution. In robot learning, therbligs offer a clear and structured method to visualize processes, as shown in Fig. 2.
Figure 2: Detailed Decomposition of a Robotic Task into therbligs. The sequence starts with Rest (R), where the robot remains idle. Transport Empty (TE) shows robot moving without any load. Delay (D) represents a pause, allowing for system adjustments or synchronization with other processes. The Grasp (G) phase involves robot’s precise movement to hold the object. In Transport Load (TL), robot carries the object to its intended destination. The Use (U) phase illustrates the robot performing the primary function with the object. Release (R) Load depicts the robot releasing the object.
Figure 3: Overview of the proposed TBBF. This pipeline integrates offline training and online testing stages. During offline training, human experts provide demonstrations and label robot trajectories into therbligs, which are then used to train the MGSF network. In the online testing stage, the trained MGSF network segments new tasks into Therblig-level actions. ActionREG registers these actions into new configurations, and LAP-VC is utilized for error compensation. Finally, YOLOv8 and PCA are used to match new configurations, ensuring robust performance across varied and complex scenarios.
An early contribution in this area is the work by Ahmadzadeh et al. [20], who introduced a novel approach for converting action sequences into symbolic representations. For action modalities, Chen et al. [21] introduced a novel approach that leverages sequential motion primitives through human demonstration. Utilizing a hierarchical BiLSTM classifier, they extracted an intuitive high-level knowledge called therbligs, which reconstructs motion information recorded from human-guided demonstrations, allowing for more general representations for different task decompositions. However, their work remains at a conceptual stage for simple therblig segmentation, and it lacks integration with vision modalities, resulting in limited generalization and scene understanding. For vision modalities, Dessalene et al. [22] introduced a rule-driven, compositional, and hierarchical action modeling method based on therbligs to analyze complex motions. This model features a novel hierarchical architecture comprising a Therblig model and an action model, utilizing vision as a medium for robot action segmentation. However, it lacks integration with action modalities based on human demonstrations, leading to a significant domain gap and inefficient capture of accurate task representations through images alone.
The concurrent study of multimodal approaches (action, vision, language) was introduced by Wang et al. [15]. This research utilized a novel LLM-based system to explore mode families within robot manipulation trajectories. The LLM can generate a list of actions involved in tasks such as scooping marbles from one bowl into another. For instance, the LLM might produce a sequence of verbs such as “reach,” “scoop,” “transport,” and “pour.” This capability allows the system to easily generate goals. These verbs are conceptually close to therbligs, providing a strong basis for task decomposition. However, while this approach closely approximates the therbligs concept, it still does not clearly establish this structure nor utilize it as the backbone for constructing the primary framework. As a result, the system lacks a well-defined structure that could enhance its robustness and reliability in diverse scenarios. Without a clear backbone, the integration of action, vision, and language modalities remains fragmented, leading to potential gaps in task representation and execution efficiency.
Figure 4: Detailed architecture of the MGSF network. The MGSF network integrates BiLSTM and Transformer sub-networks to capture sequential dependencies in robotic tasks. It utilizes a meta-recursive gated fusion mechanism to dynamically combine the outputs of these sub-networks. This fusion process adapts to different tasks, ensuring the effective integration of long-term dependencies.
In our research, we propose to use therbligs as the backbone of a robot intelligent system to enhance task understanding. This therblig-based backbone system represents a significant contribution towards a more structured, interpretable, and adaptable framework for robot task learning and execution. By integrating this approach with the foundation model, we can easily extract the detailed configurations of the objects. Thus, we can create a more robust and flexible model for robotic systems.
III Model framework
III-A TBBF: Explainable robot task understanding framework
The TBBF is designed to enhance the understanding and generalization of robotic tasks by breaking them down into fundamental units called therbligs. This framework provides a structured and modular approach, facilitating better generalization across different tasks and scenarios while offering a clearer and more interpretable structure for task execution.
In the offline training stage, we utilize the MGSF network to accurately segment tasks into therbligs, providing a detailed breakdown of the task into its constituent motions. During the online testing stage, we collect a one-shot demonstration of a new task, from which the MGSF network extracts high-level knowledge and transforms it into a structured format. This knowledge is encoded into visual data using ActionREG, integrating therbligs with the objects’ configurations in the robot’s visual field to ensure precise action registration. By using therbligs as the backbone, our framework significantly improves data efficiency and task generalization, enabling the robot to handle a wide range of scenarios with robustness and precision. The integration of LAP-VC further ensures that any visual discrepancies are corrected in real time, providing an additional layer of accuracy in task execution. As depicted in Fig. 3, this comprehensive and structured methodology enhances the robot’s ability to adapt to new tasks by leveraging prior knowledge encoded in therbligs, thus improving the interpretability, stability, and transferability of robotic learning systems.
III-B MGSF: Effieicent therbligs segmentation network
Algorithm 1 MGSF network for segmentation
  Input: $X\in\mathbb{R}^{n\times d}$, $m\in\mathbb{N}$
  Output: $\hat{Y}\in\mathbb{R}^{n\times O}$
  Initialize: $\theta_{L},\theta_{T},\theta_{g},\theta_{m},\theta_{c}$
  Initialize meta parameter vector $M\in\mathbb{R}^{m}$
  Define the number of fusion steps $T\in\mathbb{N}$
  for $i=1$ to $n$ do
     $h_{i}^{L}\leftarrow\text{BiLSTM}(x_{i};\theta_{L})$BiLSTM hidden states}
     $h_{i}^{T}\leftarrow\text{Transformer}(x_{i};\theta_{T})$Transformer hidden states}
     $c_{i}\leftarrow[h_{i}^{L}\oplus h_{i}^{T}]$
     $F=F^{(0)}$Initial fusion output}
     for $t=1$ to $T$ do
        $\theta_{g}\leftarrow f(M;\theta_{m})$Gate parameters from meta-network}
        Compute gate values $G\in\mathbb{R}^{\dim(F)}$ using $\theta_{g}$
        $G=\sigma(\theta_{g}\cdot c_{i})$Gate values using sigmoid function}
        Update fusion output $F\leftarrow G\odot c_{i}+(1-G)\odot F$where $\odot$ denotes element-wise multiplication}
        $F=G\odot c_{i}+(1-G)\odot F$Updated fusion output}
     end for
     Compute predicted output $\hat{y}_{i}\leftarrow\text{Classifier}(F;\theta_{c})$
     $\hat{y}_{i}=\text{Classifier}(F;\theta_{c})$Predicted output}
     Append predicted output to $\hat{Y}\leftarrow\hat{Y}\cup\{\hat{y}_{i}\}$
     $\hat{Y}=\hat{Y}\cup\{\hat{y}_{i}\}$Update output set}
  end for
Our MGSF network, illustrated in Fig. 4 and detailed in Algorithm 1, advances the current state of the art in understanding robot task actions. By combining meta-learning with adaptive gated fusion within a unified framework, this model significantly enhances robots’ ability to comprehend and execute sequential actions across various environments. Inspired by MetaGross [23], our MGSF network incorporates meta-gating and recursive parameterization in a recurrent model. However, MetaGross lacks a dedicated fusion process and struggles to integrate different aspects of the input data effectively, limiting its ability to leverage diverse features and achieve robust performance.
Figure 5: Details of the action registration and context matching process. This figure illustrates the stages of action registration from initial scenarios to new configurations. Key therbligs such as ’Grasp’ and ’Use’ are identified from expert demonstrations using the SAM model and matched to object masks and trajectories. The therblig-based YOLOv8 model is employed for task-related object matching, while inverse kinematics ensures accurate positioning.
To address these limitations, our MGSF model introduces a dynamic hybrid architecture that combines the strengths of both BiLSTM and Transformer sub-networks with a novel adaptive gated fusion mechanism. This architecture features a meta-recursive gated fusion unit that dynamically adapts to integrate model outputs, thereby enhancing performance across diverse tasks. Unlike the static gating in MetaGross, our adaptive gated fusion mechanism allows for more flexible and responsive integration of sequential data, ensuring that long-term dependencies are effectively captured and processed. By leveraging the strengths of both BiLSTM and Transformer sub-networks, the MGSF network excels in handling complex sequences with greater precision. The meta-learning component dynamically adjusts to the changing context of tasks, ensuring that the model remains accurate and applicable across different situations. This fusion process, informed by meta-learning, allows the MGSF model to provide a thorough analysis of each action, thereby improving the robot’s ability to perform a wide range of tasks with enhanced recall and adaptability.
III-C ActionREG: SAM-driven action registration network
A cornerstone of our TBBF is the ActionREG, designed for reasoning about and regressing object configurations (Fig. 5). This innovative component integrates therbligs’ prior knowledge into the SAM model as prompts, enabling sophisticated reasoning about the objects involved in robotic tasks. The integration of this knowledge serves as a guide for the model, improving its ability to understand and predict the configurations of objects within a task-specific context.
Algorithm 2 ActionREG and Context Matching
0:  Online demonstration $\mathbf{D}$, Therbligs segmentation model $M_{\text{MGSF}}$, Hand-eye calibration matrix $\mathbf{H}$, SAM model $M_{\text{SAM}}$, YOLOv8 model $M_{\text{YOLO}}$, Background prior $\mathbf{B}$, New environment image $\mathbf{I}_{\text{new}}$, Reference image $\mathbf{I}_{\text{ref}}$
0:  Object positions and orientations $\mathbf{P}_{\text{new}},\mathbf{O}_{\text{new}}$
  $\mathbf{S}_{\text{therbligs}}=M_{\text{MGSF}}(\mathbf{D})$
[…]
  $\mathbf{M}=\bigcup_{k\in\mathbf{K}}M_{\text{SAM}}(\mathbf{H}\cdot\mathcal{G}(\mathbf{S}_{\text{therbligs}},k),\mathbf{B})$
  for each $\mathbf{m}_{k}\in\mathbf{M}$ do
     $\mathbf{Box}_{k}=M_{\text{YOLO}}(\mathbf{I}_{\text{new}},\text{ComputeArea}(\mathbf{m}_{k}))$
     $\mathbf{F}_{\text{new}}=\text{SIFT}(\mathbf{Box}_{k})$, $\mathbf{F}_{\text{ref}}=\text{SIFT}(\mathbf{I}_{\text{ref}})$
     $\mathbf{M}_{\text{match}}=\text{FLANN}(\mathbf{F}_{\text{new}},\mathbf{F}_{\text{ref}})$
     $\mathbf{P}_{\text{new},k}=\text{ComputePosition}(\mathbf{M}_{\text{match}})$
     $\mathbf{O}_{\text{new},k}=\text{PCA}(\mathbf{M}_{\text{match}})$
  end for
  return  $\mathbf{P}_{\text{new}},\mathbf{O}_{\text{new}}$
Through ActionREG, we can easily extract the task-related object mask and the operational workspace mask. Utilizing a hierarchical matching system based on SAM (Segment Anything Model), YOLOv8 (You only look once version 8), SIFT (Scale-Invariant Feature Transform), FLANN (Fast Library for Approximate Nearest Neighbors), and PCA (Principal Component Analysis), we can identify task-related objects and determine their configurations in new scenarios. Specifically, $\mathbf{M}=\bigcup_{k\in\mathbf{K}}M_{\text{SAM}}(\mathbf{H}\cdot\mathcal{G}(\mathbf{S}_{\text{therbligs}},k),\mathbf{B})$ represents object mask segmentation using the SAM model, while $\mathbf{Box}_{k}=M_{\text{YOLO}}(\mathbf{I}_{\text{new}},\text{ComputeArea}(\mathbf{m}_{k}))$ denotes bounding box detection through YOLOv8. The feature extraction is handled by $\mathbf{F}_{\text{new}}=\text{SIFT}(\mathbf{Box}_{k})$ using SIFT, feature matching is performed by $\mathbf{M}_{\text{match}}=\text{FLANN}(\mathbf{F}_{\text{new}},\mathbf{F}_{\text{ref}})$ using FLANN, and orientation extraction is achieved via $\mathbf{O}_{\text{new},k}=\text{PCA}(\mathbf{M}_{\text{match}})$ using PCA.
ActionREG excels in performing one-shot regression of object configurations, meaning it can accurately predict an object’s configuration in new and unseen environments using just a single example or image. This capability is vital for robotic systems operating in dynamic, real-world settings, ensuring robust and adaptable performance.
III-D LAP-VC: LLM-Alignment Policy for Visual Correction
Expert demonstration may have some errors: robot end-effector grasping point may not always be perpendicular to the object itself. In addition, hand-eye calibration may also introduce some errors. These issues in real-world experiments may cause incorrect position estimation and affect our action registration results, especially in the grasp stage. To address these challenges, we introduced a novel method (Fig. 6) using LLM to correct these errors. By inputting our predicted points along with the scenario image into the LLM and utilizing a pre-built prompt, the LLM outputs corrected points. This method significantly reduces the impact of poor demonstrations and cumulative system errors, minimizing the need for highly accurate expert demonstrations.
Figure 6: Application of LAP-VC in robotic tasks. This process involves inputting predicted points along with the scenario image into the LLM. The pre-built prompt guides the LLM to output corrected points, compensating for errors due to factors such as imperfect grasping angles or hand-eye calibration inaccuracies.
IV Experiments and Setup
To demonstrate the model’s generality, we developed the Omniverse Robot System (ORS) for convenient data collection. Offline data collection involved two individuals: one performing expert demonstrations and the other labeling the robot task’s status. The system collected data on robot joint angles, end effector pose, joint speed, force, and torque readings from the ’Robotiq’ sensor, task images, and real-time labels. We gathered data from six tasks (tool-pick&place, crossbeam-cutting, bricks-gluing, tissue-sweeping, surface-wiping, cup-pouring) with expert demonstrations. For online testing, a human expert performed new tasks in unseen scenarios, with ORS recording robot states and scenario images before and after the task. We tested five challenging tasks (board-rolling, foamblock-flipping, plate-scrubbing, spoon-tilting, and paper-stamping) to evaluate performance.
We utilized a UR5 robotic arm with an origami gripper, inspired by Liu et al. [24], for manipulation tasks. This setup improves adaptability and precision. Visual and force feedback were captured using a RealSense D435i camera and a RoboTiq force sensor. The setup ran on Ubuntu 22.04, with ROS2 Humble for component communication and MoveIt2 for motion planning. Development was conducted in VSCode and model training utilized an NVIDIA RTX 4090 GPU.
V Results and Analysis
For offline training, we used robot states as input to segment therbligs. In this time-series segmentation task, our MGSF network outperforms other state-of-the-art methods. As shown in Table II, our network achieved the highest recall of 94.37%, surpassing traditional methods like Cascade SVM and deep learning models such as CNN-RNN and BiLSTM-Type3. In addition, we conducted an ablation study for our MGSF network to evaluate the impact of different components. The baseline recall of our backbone model (without fusion) was approximately 79. 77% with a transformer network. When we introduced the fusion mechanism, the recall improved to 84.29%. The addition of gate fusion mechanism boosted the recall to 90.92%. The final MGSF model, which incorporates gated fusion and meta-learning, achieved an recall of 94.37% (Table II).
TABLE II: Therblig Segmentation General Performance
Benchmark Model BCE-Loss $\downarrow$ Precision $\uparrow$ Recall $\uparrow$ F1-Score $\uparrow$ Kappa $\uparrow$ TP-Range $\uparrow$
TCNs [25] $0.623\pm 0.002$ $85.04\pm 0.79$ $88.48\pm 0.79$ $86.51\pm 0.80$ $85.54\pm 0.99$ $[85.45,92.97]$
ABLG-CNN [26] $0.472\pm 0.014$ $25.32\pm 0.55$ $24.08\pm 0.65$ $21.18\pm 0.98$ $8.23\pm 0.48$ $[18.71,32.05]$
MS-CRN [27] $0.058\pm 0.008$ $92.57\pm 1.20$ $92.55\pm 1.19$ $92.51\pm 1.21$ $90.76\pm 1.48$ $[75.58,90.11]$
BiLSTM-T3 [21] $0.137\pm 0.002$ $78.84\pm 0.56$ $81.56\pm 0.53$ $79.74\pm 0.51$ $76.82\pm 0.64$ $[73.64,86.56]$
GATv2 [28] $0.145\pm 0.008$ $81.07\pm 0.79$ $82.22\pm 0.63$ $80.47\pm 0.65$ $77.58\pm 0.78$ $[91.88,93.80]$
Reformer [29] $0.145\pm 0.005$ $77.87\pm 0.69$ $80.67\pm 0.74$ $78.79\pm 0.72$ $75.75\pm 0.91$ $[68.50,88.17]$
TFT [30] $0.307\pm 0.004$ $69.37\pm 0.46$ $71.81\pm 0.44$ $69.28\pm 0.41$ $64.45\pm 0.52$ $[62.36,82.02]$
TSMixer  [31] $0.125\pm 0.005$ $81.14\pm 0.56$ $83.51\pm 0.44$ $81.68\pm 0.46$ $79.25\pm 0.56$ $[80.06,90.30]$
Ablation Model (descending) BCE-Loss $\downarrow$ Precision $\uparrow$ Recall $\uparrow$ F1-Score $\uparrow$ Kappa $\uparrow$ TP-Range $\uparrow$
MGFN (Ours) $0.043\pm 0.007$ $94.36\pm 0.60$ $94.37\pm 0.59$ $94.36\pm 0.60$ $93.03\pm 0.73$ $[93.88,97.17]$
GrNT (no meta) [32] $0.068\pm 0.008$ $90.96\pm 0.83$ $90.92\pm 0.80$ $90.84\pm 0.83$ $88.72\pm 0.99$ $[88.74,94.73]$
Adaptive-DF (no gate) [33] $0.118\pm 0.002$ $81.86\pm 0.29$ $84.29\pm 0.24$ $82.44\pm 0.25$ $80.24\pm 0.30$ $[81.11,89.82]$
Backbone (no fusion) [34] $0.145\pm 0.004$ $77.00\pm 0.92$ $79.77\pm 0.80$ $77.94\pm 0.82$ $74.57\pm 0.99$ $[74.44,87.67]$
TABLE III: Robot Task Success Rate Comparison (Long-horizon)
Model / Task Board-rolling FoamBlock-flipping Plate-scrubbing Spoon-tilting Paper-stamping Total
SM(ST)-SimScenario 13/50 (26%) 2/50 (4%) 11/50 (22%) 0/50 (0%) 15/50 (30%) 13.7%
BC(ST)-SimScenario 6/50 (12%) 0/50 (0%) 0/50 (0%) 2/50 (4%) 0/50 (0%) 2.7%
BC(MT)-SimScenario 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0%
Ours-SimScenario 48/50 (96%) 47/50 (94%) 46/50 (92%) 48/50 (96%) 47/50 (94%) 94.4%
SM(ST)-ComScenario 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0%
BC(ST)-ComScenario 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0%
BC(MT)-ComScenario 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0/50 (0%) 0%
Ours-ComScenario 43/50 (86%) 42/50 (84%) 40/50 (80%) 34/50 (68%) 41/50 (82%) 80%
Moreover, we also analyzed the recall of different therbligs in terms of various tasks (Fig. 7). Generally, surface wiping achieved the highest segmentation results (97.17%) with six diverse robot tasks, while tissue sweeping achieved the lowest results (93.88%). This discrepancy may be attributed to the complexity of tissue sweeping actions required to effectively put tissue into the dustpan. TLoad and Release achieved the lowest recall results, around 85.56% and 83.41% respectively. However, it is important to note that even though the release recall is relatively lower, this does not significantly affect the overall results. This is because, after intricate manipulation operations, the force-torque sensors may be affected by drift.
Figure 7: Therblig segmentation recall for different robot tasks. It shows the recall rates for various therbligs across multiple robotic tasks, highlighting the performance of our segmentation network. For picka and place task, data points represented with an underscore indicate that the data is none with use.
Furthermore, as shown in Fig. 8, the LAP-VC system consistently achieves high alignment performance scores, averaging 0.896 across various tasks, outperforming traditional methods such as KNN, SIFT, ORB, AKAZE, FAST, and BRISK. Its performance is only slightly lower than that of human experts manually performing alignment. Specifically, for the Stamp task, the LAP-VC system achieved a score of 0.84. This lower score is attributed to the relatively small size of the stamp. In contrast, the Sponge task achieved a higher score of 0.94 due to the sponge’s relatively simple and uniform structure.
For the results of task execution, Table III showed that our system can achieve promising performance in terms of simple scenario (SimScenario) and complex scenario (ComScenario) with different new tasks.
Figure 8: Performance comparison across various methods and tasks.
For our SimScenario, we only consider two task-related and unseen objects appearing in our vision area. For ComScenario, we add 3-6 unrelated and unseen objects together with our task-related objects to mimic the real-world cluttered environments. For SimScenario mode, our system achieved around 94.4% average success rate for five tasks, while our success rate decreased to 80% if we switched to ComScenario mode. We find that spoon-tilting has a lower success rate because it requires a more dynamic and complex trajectory while robot solvers easily get trapped by singularity. We compared our system against Single-Task State Machine (ST-SM), Single-Task Behavior Clone (ST-BC), and Multi-Task Behavior Clone (MT-BC) baselines. State machines, built for single tasks, use pre-built policies and vision features to generate trajectories. In SimScenario, state machines achieve some success with well-built policies and excellent vision features. However, different tasks require different policies, significantly increasing system complexity, as noted by Luo et al. [16]. We collected 250 expert demonstrations for BC training and found that MT-BC training is challenging with limited data. ST-BC showed converging loss but resulted in unstable trajectories and task failures with minor errors. Applying BC to ComScenario led to divergent training and unpredictable trajectories.
ACT and diffusion policies were not used as baselines for a few reasons. Our system performs one-shot tasks and operates in zero-shot scenarios without requiring extensive pre-training for scene objects, thus needing less data. In contrast, ACT approaches typically require large datasets. Additionally, our system handles long-horizon tasks involving comprehensive action sequences, while diffusion policies focus on discrete sub-tasks. Therefore, a direct comparison may not provide meaningful insights. This distinction highlights our system’s data efficiency and generalization.
VI Conclusion and Future Work
In this paper, we presented a novel TBBF to enhance the understanding and execution of robotic tasks. By decomposing complex tasks into fundamental therbligs, our innovative approach provides a structured and interpretable representation of robotic actions. The MGSF network was employed for accurate therblig segmentation during the offline training stage, while ActionREG efficiently facilitated precise action registration and integration of therbligs with object configurations during the online testing stage.
Our experimental results demonstrate the effectiveness of our framework, showcasing high recall in therblig segmentation and robust performance in real-world robot task execution. Specifically, we achieved results with 94.37% recall in therblig segmentation and impressive successful execution rates of 94.4% for new and long-horizon tasks in simple scenarios, and 80% in complex scenarios. However, some limitations should be addressed in future work. Firstly, our model relies heavily on standard expert demonstrations, which must follow specific protocols. It struggles with extremely noisy or unprofessional demonstrations. Secondly, our model, like many existing approaches, is affected by object shadows and light reflections. Additionally, we focused on the generation of 2D object configurations and did not account for 3D object configurations, which could be improved with depth information.
Future research will focus on refining the TBBF to enhance its robustness and adaptability in more complex and dynamic environments. We will also incorporate more user data to train a more robust action segmentation network. Furthermore, we plan to extend the applicability of our methods to a broader range of robotics platforms, facilitating the transfer of robot knowledge from the UR5 to other platforms such as Panda or Kinova robots.